OpenStudy (anonymous):

Find the slope of the line whose equation is 2x-5y=7

OpenStudy (ilovecake):

ok any ideas or answer choices?

OpenStudy (texaschic101):

put the equation in y = mx + b form, and the number in the m position is your slope

OpenStudy (texaschic101):

step 1 : subtract 2x from both sides step 2 : divide both sides by -5 your equation will then be in y = mx + b form....now look at the number in the m position...that is your slope

OpenStudy (anonymous):

I don't have any choices

OpenStudy (anonymous):

Okay

OpenStudy (anonymous):

Is that all ?

OpenStudy (texaschic101):

yes....because in y = mx + b form, the number in the m position (the number right before the x) is the slope

OpenStudy (anonymous):

Okay thank you

OpenStudy (texaschic101):

do you still need help

OpenStudy (anonymous):

No I'm good

OpenStudy (texaschic101):

good to hear....if you are still having problems, just tag me :)

OpenStudy (ilovecake):

@texaschic101 She is right

OpenStudy (ilovecake):

@texaschic101 I mean you are right